OpenStudy (anonymous):

Subtract the complex numbers: (7+4i)-(5-6i) Is the answer 2-2i??

OpenStudy (sidsiddhartha):

(7+4i)-(5-6i)=7+4i-5+6i=2+10i=2(1+5i)

OpenStudy (anonymous):

sadly no... your imaginary term is wrong. Remember you are subtracting a negative number, so it becomes positive

OpenStudy (anonymous):

2 + 10i is the right answer

OpenStudy (anonymous):

Oh okay!! So like sid said, it is 2+10i. Got it now!

OpenStudy (sidsiddhartha):

yup